OpenStudy (anonymous):

Simplify. (6.25x10-17)-2 Use scientific notation. Use the multiplication symbol in the math palette as needed. The -17 and the -2 are exponents

OpenStudy (anonymous):

(6.25*10-17)-2 Multiply 6.25 by 10 to get 62.5. (62.5-17)-2 Subtract 17 from 62.5 to get 45.5. (45.5)-2 Remove the parentheses around the expression 45.5. 45.5-2 Subtract 2 from 45.5 to get 43.5. 43.5

OpenStudy (anonymous):

that help out

OpenStudy (anonymous):

the -2 and the -17 are exponents

OpenStudy (anonymous):

\[(6.25 \times 10^{-17})^{-2}=6.25^{-2} \times 10 ^{-17 \times -2}=\frac{1}{6.25^2}\times 10^{34}\]\[=\frac{1}{(\frac{25}{4})^2} \times 10^{34}= \frac{16}{625}\times 10^{34}\]

OpenStudy (anonymous):

\[=2.56 \times 10^{32}\]to 3 sig. fig.

OpenStudy (anonymous):

toodles, is that okay?